Find the interval(s) where the function is increasing and the interval(s) where it is decreasing. (Enter your answers using interval notation. If an answer cannot be expressed as an interval, enter EMPTY or ∅.)Find the interval(s) where the function is increasing and the interval(s) where it is decreasing. (Enter your answers using i
Answer #1

Given f(x) = 4x + 100²-968-9 fkucar 3x²+20x-96 fkasso x²+200-96=0 +5% - 24-0 --3)(x+8)=0 21 = 3,-8 2 a + t 3 Constructing Sig